OpenStudy (anonymous):

Solve the equation. 3x - 10 = 3(x - 4) + 2 a. 8 b. 0 c. infinite solutions d. no solution

OpenStudy (ivettef365):

infinite solution because 3x-10 = 3x-12=2 3x-10 = 3x-10 therefore any number will make true the equation

OpenStudy (ivettef365):

any number for x

OpenStudy (anonymous):

Thank you I really need these.